The refractive index of the core and cladding materials of an optical fibre is 1.46

and 1.38, respectively. Calculate the critical angle, numerical aperture and

acceptance angle at the air-fibre interface.


1
Expert's answer
2021-05-03T10:29:55-0400

Answer

Critical angle

"\theta=sin^{-1}(\frac{1.38}{1.46}) =70.9°"


Numerical aperture

"NA=(1.46^2-1.38^2) ^{0.5}=0.476"


Acceptance angle



"\theta=sin^{-1}(0.476) =28.46°"